OpenStudy (anonymous):

The system of equations is coincident. What are the missing values? -2x+5y= -12 -6+ y=

OpenStudy (anonymous):

@tiff9702

OpenStudy (anonymous):

hay tiff

OpenStudy (tiff9702):

hi

OpenStudy (anonymous):

can u help me on this. this question as me stomped

OpenStudy (tiff9702):

i'm working on it

OpenStudy (anonymous):

k

OpenStudy (tiff9702):

-2x + 5y = -12 -6x + _ y = _ --------------- If the lines are to be coincident, then they will coincide, be the same line. So, the equations have to be equivalent. Looking at the x terms, 3 times -2x = -6x So, multiply 3 times 5 to get 15 for the missing coefficient of y. Multiply -12 by 3 to get the missing constant term.

OpenStudy (anonymous):

so i got -36 and 15 right?

OpenStudy (tiff9702):

yep

OpenStudy (anonymous):

ok i get now xD thx for helping

OpenStudy (tiff9702):

np